# FoREST: Frame of Reference Evaluation in Spatial Reasoning Tasks

### v2

@@ Abstract
Spatial reasoning is a fundamental aspect of human intelligence. One key concept in spatial cognition is the Frame of Reference (FoR), which identifies the perspective of spatial expressions. Despite its significance, FoR has received limited attention in AI models that need spatial intelligence. There is a lack of dedicated benchmarks and in-depth evaluation of large language models (LLMs) in this area. To address this issue, we introduce the Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST) benchmark, designed to assess FoR comprehension in LLMs. We evaluate LLMs on answering questions that require FoR comprehension and layout generation in text-to-image models using FoREST. Our results reveal a notable performance gap across different FoR classes in various LLMs, affecting their ability to generate accurate layouts for text-to-image generation. This highlights critical shortcomings in FoR comprehension. To improve FoR understanding, we propose Spatial-Guided prompting, which improves LLMs’ ability to extract essential spatial concepts. Our proposed method improves overall performance across spatial reasoning tasks.
@@
Tanawan Premsri Department of Computer Science Michigan State University premsrit@msu.edu                        Parisa Kordjamshidi Department of Computer Science Michigan State University kordjams@msu.edu
@@ 1 Introduction
Spatial reasoning plays a significant role in human cognition and daily activities. It is also a crucial aspect in many AI problems, including language grounding (Zhang and Kordjamshidi, 2022; Yang et al., 2024), navigation (Yamada et al., 2024), computer vision (Liu et al., 2023; Chen et al., 2024), medical domain (Gong et al., 2024), and image generation (Gokhale et al., 2023). One key concept in spatial reasoning is the Frame of Reference (FoR), which identifies the perspective of spatial expressions. FoR has been studied extensively in cognitive linguistics (Edmonds-Wathen, 2012; Vukovic and Williams, 2015). Levinson (2003) initially defines three FoR classes: relative, based on the observer’s perspective; intrinsic, based on an inherent feature of the reference object; and absolute, using environmental cues like cardinal directions -See Figure 1. This framework was expanded by  Tenbrink (2011) to create a more comprehensive framework, serving as the basis of our work. Understanding FoR is important for many applications, especially in embodied AI. In such applications, an agent must simultaneously comprehend multiple perspectives, including the one from the instruction giver and from the instruction follower, to communicate and perform tasks effectively. However, recent spatial evaluation benchmarks have largely overlooked FoR. For example, the text-based benchmarks Shi et al. (2022); Mirzaee and Kordjamshidi (2022); Rizvi et al. (2024) and text-to-images benchmarks (Gokhale et al., 2023; Huang et al., 2023; Cho et al., 2023a, b) assume a fixed perspective for all spatial expressions. This inherent bias limits situated spatial reasoning, restricting adaptability in interactive environments where perspectives can change.
@@
Figure 1: Illustration of FoR classes. The cat is the locatum, the car is the relatum, and the arrow denotes the perspective.
@@
To systematically investigate the role of FoR in spatial understanding and create a new resource, that is, Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST), FoREST is designed to evaluate models’ ability to comprehend FoR from textual descriptions and extend this evaluation to grounding and visualization. Our benchmark includes spatial expressions with FoR ambiguity—where multiple FoRs may apply to the described situation—and spatial expressions with only a single valid FoR. This design allows evaluation of the models’ understanding of FoR in both scenarios. We evaluate several LLMs in a QA setting that require FoR understanding and apply the FoR concept in text-to-image models. Our findings reveal performance differences across FoR classes and show that LLMs exhibit bias toward specific FoRs when handling ambiguous cases. This bias extends to layout-diffusion models, which rely on LLM-generated layouts in the image generation pipeline. To enhance FoR comprehension in LLMs, we propose Spatial-Guided prompting, which enables models to analyze and extract additional spatial information, including directional, topological, and distance relations. We demonstrate that incorporating spatial information improves question-answering and layout generation, ultimately enhancing text-to-image generation performance.
@@
Our contribution111code and dataset available at GitHub repository. are summarized as follows, 1. We introduce the FoREST benchmark to evaluate LLMs’ FoR comprehension in a QA setting systematically. 2. We analyze the impact of FoR information on text-to-image generation using multiple diffusion models. 3. We propose a prompting approach that generates spatial information, which can be incorporated into QA and layout diffusion to enhance performance.
@@ 2 Spatial Primitives
Spatial Relations. When dealing with spatial knowledge representation and reasoning, three main relations are often considered, directional, topological, and distance (Hernández, 1994; Cohn and Renz, 2008; Kordjamshidi et al., 2011). Directional describes an object’s direction based on specific coordinates, e.g., left and right. Topological describes the containment between two objects, such as inside. Distance describes qualitative and quantitative relations between entities. Examples of qualitative are far, and quantitative are 3km.
Spatial Frame of Reference. We use four frames of references investigated in the cognitive linguistic studies Tenbrink (2011). These are defined based on the concept of Perspective, which is the origin of a coordinate system to determine the direction. The four frames of reference are defined as follows.
1. External Intrinsic describes a spatial relation from the relatum’s perspective, where the relatum does not contain the locatum. The top-right image in Figure 1 illustrates this with the sentence, A cat is to the right of the car from the car’s perspective.
@@
3. Internal Intrinsic describes a spatial relation from the relatum’s perspective, where the relatum contains the locatum. The bottom-right image in Figure 1 shows this with the sentence, A cat is inside and back of the car from the car’s perspective.
4. Internal Relative describes a spatial relation from the observer’s perspective where the locatum is inside the relatum. The bottom-left image in Figure 1 shows this FoR with the sentence, A cat is inside and to the left of the car from my perspective.
Figure 2: Pipeline for dataset creation, starting from selecting a locatum and relatum from available objects and then applying a spatial template to generate the spatial expression ($T$). FoRs are assigned based on relatum properties. $T$ is categorized based on the number of FoRs. For example, A cat is to the right of a dog (with two possible FoRs: external intrinsic and external relative) belongs to the A-split. Then, its disambiguated version (A cat is to the right of a dog from the dog’s perspective) is added to the C-split. Next, if applicable, a relatum’s orientation is included for visualization and question generation. Finally, Unity3D generates scene configurations, and question-answer pairs are derived from $T$.
@@ 3 FoREST Dataset Construction
To systematically evaluate LLM on the frame of reference (FoR) recognition, we introduce the Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST) benchmark. Each instance in FoREST consists of a spatial context ($T$), a set of corresponding FoR ($FoR$) which is a subset of {external relative, external intrinsic, internal intrinsic, internal relative}, a set of questions and answers ($\{Q,A\}$), and a set of visualizations ($\{I\}$). An example of $T$ is A cat is to the right of a dog. A dog is facing toward the camera. The FoR of $T$ is {external intrinsic, external relative}. A question-answer is $Q$ = Based on the camera’s perspective, where is the cat from the dog’s position?, $A$ = {left, right}. There is an ambiguity in the FoR for this expression. Thus, the answer will be left if the model assumes the external relative. Conversely, it will be right if the model assumes the external intrinsic. The visualization of this example is in Figure 2.
@@ 3.1 Context Generation
We select two distinct objects—a relatum ($R$) and a locatum ($L$)—from a set of 20 objects and apply them to a Spatial Relation template, <$L$> <spatial relation> <$R$> to generate the context $T$. FoRs for $T$ are determined based on the properties of the selected objects. Depending on the number of possible FoRs, $T$ is categorized as ambiguous (A-split), where multiple FoRs apply, or clear (C-split), where only one FoR is valid. We further augment the C-split with disambiguated spatial expressions derived from the A-split, as shown in Figure 2.
@@ 3.2 Categories based on Relatum Properties
Using the FoR classes in Section 2, we identify two key properties contributing to FoR ambiguity. The first property is the relatum’s intrinsic direction. It creates ambiguity between intrinsic and relative FoRs as spatial relations may originate from either relatum’s or observer’s perspectives. The second is the relatum’s affordance as a container. It introduces the ambiguity between internal and external FoR, as spatial relations may refer to either inside or outside of the relatum. Note that containment is defined as the ability of the relatum to contain the locatum, considering both objects’ sizes. Based on these properties, we define four distinct cases: Cow, Box, Car, and Pen.
@@
Case 1: Cow Case. In this case, the selected relatum has intrinsic directions but cannot be the container for the locatum. An example is a cow. In such cases, the relatum potentially provides a perspective for spatial relations. The applicable FoR classes are $FoR$ = {external intrinsic, external relative}. We augment the C-split with expressions of this case but include the perspective to resolve their ambiguity. To specify the perspective, we use predefined templates for augmenting clauses, such as from {relatum}’s perspective for external intrinsic or from the camera’s perspective for external relative. For example, if the context is A cat is to the right of the cow, in the A-split. The counterparts included in the C-split are A cat is to the right of the cow from cow’s perspective. for external intrinsic and A cat is to the right of the cow from my perspective. for external intrinsic.
@@
Case 2: Box Case. The relatum in this case has the property of being a container but lacks intrinsic directions, making the internal FoR applicable. An example is a box. The applicable FoR classes are $FoR$ = {external relative, internal relative}. To include their unambiguous counterparts in the C-split, we specify the topological relation by adding inside for internal relative and outside for external relative. For example, for the sentence A cat is to the right of the box., the unambiguous $T$ with internal relative FoR is A cat is inside and to the right of the box. The counterpart for external relative is A cat is outside and to the right of the box.
@@
Case 3: Car Case. A relatum with an intrinsic direction and container affordance falls into this case, allowing all FoR classes. An obvious example is a car that can be a container with intrinsic directions. The applicable FoR classes are $FoR$ = { external relative, external intrinsic, internal intrinsic, internal relative}. To augment C-split with this case’s disambiguated counterparts, we add perspective and topology information similar to the Cow and Box cases. An example expression for this case is A person is in front of the car. The four disambiguated counterparts to include in the C-split are A person is outside and in front of the car from the car itself. for external intrinsic, A person is outside and in front of the car from the observer. for external relative, A person is inside and in front of the car from the car itself. for internal intrinsic, and A person is inside and in front of the car from the observer. for internal relative.
@@
Case 4: Pen Case. In this case, the relatum lacks both the intrinsic direction and the affordance as a container. An obvious example is a pen. Lacking these two properties, the created context has only one applicable FoR, $FoR$ = {external relative}. Therefore, we can categorize this case into both splits without any modification. An example of such a context is The book is to the left of a pen.
@@ 3.3 Context Visualization
In our visualization, complexity arises when the relatum has an intrinsic direction within the intrinsic FoR, as its orientation can complicate the spatial representation. For example, for visualization of A cat is to the right of a dog from the dog’s view, the cat can be placed in different coordinates based on the dog’s orientation. To address this issue, we add a template sentence for each direction, such as <relatum> is facing toward the camera, to specify the relatum’s orientation of all applicable $T$ for visualization and QA. For instance, A cat is to the left of a dog. becomes A cat is to the left of a dog. The dog is facing toward the camera. To avoid occlusion issues, we generate visualizations only for external FoRs, as one object may become invisible in internal FoR classes. We use only expressions in C-split since those have a unique FoR interpretation for visualization. We then create a scene configuration by applying a predefined template, as illustrated in Figure 2. Images are generated using the Unity 3D simulator Juliani et al. (2020), producing four variations per expression $T$ with different backgrounds and object positions. Further details on the simulation process are in Appendix A.2.
@@ 3.4 Question-Answering Generation
We generate questions for all generated spatial expressions ($T$). Note that we include the relatum orientation for cases where the relatum has an intrinsic direction, as mentioned in the visualization. Our benchmark includes two types of questions. The first type asks for the spatial relation between two given objects from the camera’s perspective, following predefined templates such as, Based on the camera’s perspective, where is the locatum relative to the relatum’s position? Template variations are made based on GPT4o; details are provided in Appendix A.2.4. The second type of question queries the spatial relation from the relatum’s perspective. This question type follows the same templates but replaces the camera with the relatum. The first type of question is generated for all $T$, while the second type is only generated for $T$ where the relatum has intrinsic direction and a perspective can be defined accordingly. Answers are determined based on the corresponding FoRs, the spatial relation in $T$, and the relatum’s orientation when applicable.

### v3

@@ Abstract
Spatial reasoning is a fundamental aspect of human intelligence. One key concept in spatial cognition is the Frame of Reference (FoR), which identifies the perspective of spatial expressions. Despite its significance, FoR has received limited attention in AI models that need spatial intelligence. There is a lack of dedicated benchmarks and in-depth evaluation of large language models (LLMs) in this area. To address this issue, we introduce the Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST) benchmark, designed to assess FoR comprehension in LLMs. We evaluate LLMs on answering questions that require FoR comprehension and layout generation in text-to-image models using FoREST. Our results reveal a notable performance gap across different FoR classes in various LLMs, affecting their ability to generate accurate layouts for text-to-image generation. This highlights critical shortcomings in FoR comprehension. To improve FoR understanding, we propose Spatial-Guided prompting, which improves LLMs’ ability to extract primitive spatial concepts and relations. Our proposed method improves overall performance across spatial reasoning tasks.
@@
Tanawan Premsri  and Parisa Kordjamshidi Department of Computer Science and Engineering Michigan State University {premsrit, kordjams}@msu.edu
@@ 1 Introduction
Spatial reasoning plays a significant role in human cognition and daily activities. It is also a crucial aspect in many AI problems Kordjamshidi et al. (2025), including language grounding (Zhang and Kordjamshidi, 2022; Yang et al., 2024), navigation (Zhang and Kordjamshidi, 2023; Yamada et al., 2024), computer vision (Liu et al., 2023; Chen et al., 2024), medical domain (Gong et al., 2024), and image generation (Cho et al., 2023b; Gokhale et al., 2023). One key concept in spatial reasoning is the Frame of Reference (FoR), which identifies the perspective of spatial expressions. FoR has been studied extensively in cognitive linguistics (Edmonds-Wathen, 2012; Vukovic and Williams, 2015). Levinson (2003) initially defines three FoR classes: relative, based on the observer’s perspective; intrinsic, based on an inherent feature of the reference object; and absolute, using environmental cues like cardinal directions (see Figure 1). This framework was expanded by Tenbrink (2011) to a more comprehensive framework, serving as the basis of this paper. Understanding FoR is important for many applications, especially in embodied AI. In such applications, an agent must simultaneously comprehend multiple perspectives, including the one from the instruction giver and from the instruction follower, to communicate and perform tasks effectively Liu et al. (2013). However, recent spatial evaluation benchmarks have largely overlooked FoR. For example, the text-based benchmarks Mirzaee et al. (2021); Shi et al. (2022); Mirzaee and Kordjamshidi (2022); Rizvi et al. (2024) and text-to-images benchmarks (Gokhale et al., 2023; Huang et al., 2023; Cho et al., 2023a, b) assume a fixed perspective for all spatial expressions. This inherent bias limits situated spatial reasoning, restricting adaptability in interactive environments where perspectives can change.
@@
Figure 1: Illustration of FoR classes. The cat is the locatum, the car is the relatum, and arrows denote the perspective.
@@
To systematically investigate the role of FoR in spatial understanding, we create a new resource, Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST), to evaluate models’ ability to comprehend FoR from textual descriptions and extend this evaluation to grounding and visualization. Our benchmark includes spatial expressions with FoR ambiguity—cases where multiple FoRs may apply to the described situation—as well as spatial expressions with only a single valid FoR. This design allows evaluation of the models’ understanding of FoR in both scenarios. We evaluate several LLMs in a QA setting that requires FoR understanding and employ the FoR concept in text-to-image models. Our findings reveal performance differences across FoR classes and show that LLMs exhibit bias toward specific FoRs when handling ambiguous cases. This bias extends to layout-diffusion models, which rely on LLM-generated layouts in the image generation pipeline. To enhance FoR comprehension in LLMs, we propose Spatial-Guided prompting, which enables models to analyze and extract additional spatial information, including directional, topological, and distance relations. We demonstrate that incorporating spatial primitives and relations improves question-answering and layout generation, ultimately enhancing text-to-image generation performance.
@@
Our contribution111code and dataset available at  Github repository. are summarized as follows, 1) We introduce the FoREST benchmark to systematically evaluate LLMs’ FoR comprehension, 2) We analyze the impact of FoR information on text-to-image generation using diffusion models, 3) We propose a prompting approach that generates spatial primitives and relations in the chain of reasoning, which enhances the performance of QA and layout diffusion models.
@@ 2 Spatial Primitives
Spatial Relations. When dealing with spatial knowledge representation and reasoning, three main relations are often considered: directional, topological, and distal (Hernández, 1994; Cohn and Renz, 2008; Kordjamshidi et al., 2011). Directional describes an object’s direction based on specific coordinates, e.g., left and right. Topological describes the containment between two objects, such as inside. Distal describes qualitative and quantitative relations regarding the distance between entities. An example of a qualitative distal relation is far, and an example of a quantitative one is $3\text{km}$.
Frame of Reference. We use four frames of reference investigated in the cognitive linguistic studies Tenbrink (2011). These are defined based on the concept of Perspective, which is the origin of a coordinate system to determine the direction. The four frames of reference are defined as follows.
1. External Intrinsic describes a spatial relation from the relatum’s perspective, where the relatum does not contain the locatum. The top-right image in Figure 1 shows this case with the sentence, A cat is to the right of the car from the car’s perspective.
@@
3. Internal Intrinsic describes a spatial relation from the relatum’s perspective, where the relatum contains the locatum. The bottom-right of Figure 1 shows this case with the sentence, A cat is inside and back of the car from the car’s perspective.
4. Internal Relative describes a spatial relation from the observer’s perspective where the locatum is inside the relatum. The bottom-left image in Figure 1 shows this case with the sentence, A cat is inside and to the left of the car from my perspective.
Figure 2: The dataset creation pipeline. It begins by selecting a locatum and a relatum from a pre-defined list of objects and then applies templates to generate the spatial expressions ($T$). FoRs are then assigned based on the relatum properties. $T$ is categorized based on the number of applicable FoRs. For example, A cat is to the right of a dog (with two possible FoRs: external intrinsic and external relative) belongs to the A-split. Then, its disambiguated version (A cat is to the right of a dog from the dog’s perspective) is added to the C-split. Next, if applicable, the relatum orientation is included for visualization and question generation. Finally, Unity3D generates the scene configurations, and the question-answer pairs are derived from $T$.
@@ 3 FoREST Dataset Construction
To systematically evaluate LLMs on the frame of reference (FoR) recognition, we introduce the Frame of Reference Evaluation in Spatial Reasoning Tasks (FoREST) benchmark. Each instance in FoREST consists of a spatial context ($T$), a set of corresponding FoRs ($FoR$) which is a subset of {external relative, external intrinsic, internal intrinsic, internal relative}, a set of questions and answers ($\{Q,A\}$), and a set of visualizations ($\{I\}$). An example of $T$ is “A cat is to the right of a dog. A dog is facing toward the camera.” The set of applicable FoRs for $T$ is {external intrinsic, external relative}. A question-answer pair is $Q$ = “Based on the camera’s perspective, where is the cat from the dog’s position?”, $A$ = {left, right}. There is an ambiguity in the FoR for this expression. Thus, the answer will be left if the model assumes the external relative FoR. In contrast, it will be right if the model assumes the external intrinsic FoR. The visualization of this example is shown in Figure 2. Dataset statistics are provided in Table 1.
@@ 3.1 Context Generation
We select two distinct objects—a relatum ($R$) and a locatum (LL)—from a set of 20 objects and apply them to a Spatial Relation template, <LL> <spatial relation> <RR>, to generate the context $T$. FoRs for $T$ are determined based on the properties of the selected objects. Depending on the number of possible FoRs, $T$ is categorized as ambiguous (A-split), where multiple FoRs apply, or clear (C-split), where only one FoR is valid. We further augment the C-split with disambiguated spatial expressions derived from the A-split, as shown in Figure 2.
@@ 3.2 Categories based on Relatum Properties
Using the FoR classes in Section 2, we identify two key properties contributing to FoR ambiguity. The first property is the relatum’s intrinsic direction. It creates ambiguity between intrinsic and relative FoRs, since spatial relations may originate from either the relatum’s or the observer’s perspective. The second property is the relatum’s affordance as a container. It introduces the ambiguity between internal and external FoR, as spatial relations may refer to either inside or outside of the relatum. Note that containment is defined as the ability of the relatum to contain the locatum, considering both objects’ sizes. Based on these properties, we define four distinct cases: Cow, Box, Car, and Pen.
@@
Case 1: Cow Case. In this case, the selected relatum has intrinsic directions but cannot be the container for the locatum. An example object is a cow. In such cases, the relatum provides a perspective for spatial relations. The applicable FoR classes are $FoR$ = {external intrinsic, external relative}. We augment the C-split with expressions of this case, but include the perspective to resolve their ambiguity. To specify the perspective, we use predefined templates for augmenting clauses, such as from {relatum}’s perspective for external intrinsic or from the camera’s perspective for external relative. For example, if the context is “A cat is to the left of the cow”, in the A-split, the counterparts included in the C-split are “A cat is to the left of the cow from the cow’s perspective.” for the external intrinsic and “A cat is to the left of the cow from my perspective.” for the external intrinsic.
@@
Case 2: Box Case. In this case, the relatum has the container affordance but lacks intrinsic directions, e.g, a box. The applicable FoR classes are $FoR$ = {external relative, internal relative}. To include unambiguous counterparts in the C-split, we specify a topological relation by adding inside for internal relative and outside for external relative. For example, given the sentence “A cat is to the left of the box.”, the unambiguous $T$ with internal relative FoR is “A cat is inside and to the left of the box.” The counterpart for external relative is “A cat is outside and to the left of the box.”
@@
Case 3: Car Case. A relatum with an intrinsic direction and container affordance falls into this case. An example object is a car. The applicable FoR classes are $FoR$ = { external relative, external intrinsic, internal intrinsic, internal relative}. To augment C-split with the disambiguated counterparts of such cases, we add perspective and topology information to the sentences similar to the Cow and Box cases. An example expression for this case is A person is in front of the car. The four disambiguated counterparts to include in the C-split are “A person is outside and in front of the car from the car itself.” for external intrinsic, “A person is outside and in front of the car from the observer.” for external relative, “A person is inside and in front of the car from the car itself.” for internal intrinsic, and “A person is inside and in front of the car from the observer.” for internal relative.
@@
Case 4: Pen Case. In this case, the relatum lacks both the intrinsic direction and the affordance as a container. An example object is a pen. Lacking these two properties, the created context has only one applicable FoR, $FoR$ = {external relative}. Therefore, we can categorize this case into both splits without any modification. An example of such a context is “The book is to the left of a pen.”
@@ 3.3 Context Visualization
In our visualization, complexity arises when the relatum has an intrinsic direction, as its orientation can complicate the spatial representation. For example, for the visualization of A cat is to the right of a dog from the dog’s view, the cat can be placed in different coordinates based on the dog’s orientation. To address this issue, we add a template sentence for each direction, such as <relatum> is facing toward the camera, to specify the relatum’s orientation of all applicable $T$ for visualization and QA. For instance, “A cat is to the left of a dog.” becomes “A cat is to the left of a dog. The dog is facing toward the camera.”. To avoid occlusion, we generate visualizations only for external FoRs, since one object may become invisible in internal FoR classes. We use only C-split expressions, as they have a unique FoR interpretation for visualization. A scene configuration is then created by applying a predefined template, as illustrated in Figure 2. The images are generated using the Unity 3D simulator Juliani et al. (2020), producing four variations per expression $T$ with different backgrounds and object positions. Further details of the creation process are provided in Appendix A.
@@ 3.4 Question-Answering Generation
We generate corresponding questions for each spatial expression ($T$). When the relatum has an intrinsic direction, we also include its orientation as described in Section 3.3. Our benchmark contains two question types. The first asks for the spatial relation between two objects from the camera’s perspective, following predefined templates such as, Based on the camera’s perspective, where is the locatum relative to the relatum’s position? The second queries the relation from the relatum’s perspective, using the same templates but replacing the camera with the relatum. The first type is generated for all $T$, while the second applies only when the relatum has an intrinsic direction. Answers are determined by the corresponding FoR, the spatial relation expressed in $T$, and the relatum’s orientation when applicable. Question template variations were generated using GPT-4o, with details provided in Appendix A.3. To demonstrate that the order of the locatum and the relatum in the question does not affect LLMs’ performance, we also evaluate Qwen2-72B on templates with the reversed order of the locatum and the relatum in Appendix F.2. Qwen2-72B performs similarly on both orders, so we use only the question templates where the locatum precedes the relatum in the remaining experiments.
A-split
Case Context QA-CP QA-RP T2I
Cow Case 792 3168 3168 3168
Car Case 128 512 512 512
Box Case 120 120 120 120
Pen Case 488 488 488 488
Total 1528 4288 4288 4288
C-Split
FoR Class Context QA-CP QA-RP T2I
External Relative 1528 4288 3680 4288
External Intrinsic 920 3680 3680 3680
Internal Intrinsic 128 512 512 0
Internal Relative 248 632 512 0
Total 2824 9112 8384 7968
Table 1: Dataset statistics of FoREST A-split and C-split portions. QA-CP and QA-RP are question-answer pairs with camera perspective and question-answer pairs with relatum perspective, respectively. T2I refers to the prompt used in Text-to-Image experiments.